OpenStudy (lena772):

An investor comes to your office. He says that if you give him the $5,000 he will add on an additional $50 each year to what he owes you. Create the function for this investor's plan.

OpenStudy (lena772):

@ranga

OpenStudy (anonymous):

R = x +50 -5000 R = revenue , money brought it x = years

OpenStudy (anonymous):

the x is the years, 50 each year, and you give away 5000, so you subtract that from the f-n im new to this but it seems ok to me

OpenStudy (ranga):

So he will add the same $50 each year. So for t years it will be: 50 * t The balance on the account will be: 5000 + 50*t

OpenStudy (lena772):

can I put 5000+50t

OpenStudy (ranga):

Yes. The maturity amount A after t years will be: A = 5000 + 50t

OpenStudy (lena772):

f(t)=5000+ft

OpenStudy (ranga):

not ft. but 50t

OpenStudy (lena772):

It's ok to put A for the function?

OpenStudy (ranga):

as long as each symbol is defined clearly. You can say A or even A(t) is the maturity amount after t years. If you prefer to use f that is fine too. Then you will say f(t) = 5000 + 50t where f(t) represents the amount after t years.

OpenStudy (lena772):

what about for the previous functions we did

OpenStudy (ranga):

functions dont always have to be f(x) or y = ... etc. f(x) = mx + b is a function. y = mx + b is also a function. z = 2t^2 + 3t is also a function. So it does not matter as along as you define what each term represents.

OpenStudy (lena772):

ok thank you!

OpenStudy (ranga):

alright. yw.